OpenStudy (anonymous):

PLEASE HELP WILL GIVE TONS OF POINTS AND KISSES Simplify. 4(6 + 7)

OpenStudy (anonymous):

A(B + C) = AB + AC

OpenStudy (anonymous):

um why are there leetters were working with numbers rn

OpenStudy (anonymous):

Do you know PEMDAS?

OpenStudy (anonymous):

no :(

OpenStudy (agent0smith):

4(6+7) = 4(13) = multiply them together

OpenStudy (agent0smith):

or 4*6 + 4*7